You have just been hired as a new management trainee by Earrings Unlimited, a distributor of earrings to various retail outlets located in shopping malls across the country. In the past, the company has done very little in the way of budgeting and at certain times of the year has experienced a shortage of cash. Since you are well trained in budgeting, you have decided to prepare a master budget for the upcoming second quarter. To this end, you have worked with accounting and other areas to gather the information assembled below.

The company sells many styles of earrings, but all are sold for the same price—$16 per pair. Actual sales of earrings for the last three months and budgeted sales for the next six months follow (in pairs of earrings):

January (actual) 23,400 June (budget) 53,400
February (actual) 29,400 July (budget) 33,400
March (actual) 43,400 August (budget) 31,400
April (budget) 68,400 September (budget) 28,400
May (budget) 103,400

The concentration of sales before and during May is due to Mother’s Day. Sufficient inventory should be on hand at the end of each month to supply 40% of the earrings sold in the following month.

Suppliers are paid $5.70 for a pair of earrings. One-half of a month’s purchases is paid for in the month of purchase; the other half is paid for in the following month. All sales are on credit. Only 20% of a month’s sales are collected in the month of sale. An additional 70% is collected in the following month, and the remaining 10% is collected in the second month following sale. Bad debts have been negligible.

Monthly operating expenses for the company are given below:

Variable:
Sales commissions 4 % of sales
Fixed:
Advertising $ 370,000
Rent $ 35,000
Salaries $ 140,000
Utilities $ 15,500
Insurance $ 4,700
Depreciation $ 31,000

Insurance is paid on an annual basis, in November of each year.

The company plans to purchase $24,500 in new equipment during May and $57,000 in new equipment during June; both purchases will be for cash. The company declares dividends of $27,750 each quarter, payable in the first month of the following quarter.

The company’s balance sheet as of March 31 is given below:

Assets
Cash $ 91,000
Accounts receivable ($47,040 February sales; $555,520 March sales) 602,560
Inventory 155,952
Prepaid insurance 29,500
Property and equipment (net) 1,120,000
Total assets $ 1,999,012
Liabilities and Stockholders’ Equity
Accounts payable $ 117,000
Dividends payable 27,750
Common stock 1,140,000
Retained earnings 714,262
Total liabilities and stockholders’ equity $ 1,999,012

The company maintains a minimum cash balance of $67,000. All borrowing is done at the beginning of a month; any repayments are made at the end of a month.

The company has an agreement with a bank that allows the company to borrow in increments of $1,000 at the beginning of each month. The interest rate on these loans is 1% per month and for simplicity we will assume that interest is not compounded. At the end of the quarter, the company would pay the bank all of the accumulated interest on the loan and as much of the loan as possible (in increments of $1,000), while still retaining at least $67,000 in cash.

1. a. A sales budget, by month and in total.

    b. A schedule of expected cash collections, by month and in total.

    c. A merchandise purchases budget in units and in dollars. Show the budget by month and in total.

    d. A schedule of expected cash disbursements for merchandise purchases, by month and in total.

I Only really need help with 1B. Thank you!
Answer #1

Schedule of Expected Cash Collections:

April

May

June

Quarter

February sales (29400*16=470400)*10%

47040

47040

March sales (43400*16=694400)*70%, 10%

486080

69440

555520

April sales (68400*16=1094400)*20%, 70%, 10%

218880

766080

109440

1094400

May sales (103400*16=1654400)*20%, 70%

330880

1158080

1488960

June sales (53400*16=854400)*20%

170880

170880

Total Cash Collections

752000

1166400

1438400

3356800

If you like the answer, kindly give a ?.